OpenStudy (anonymous):

The vertex of a parabola is (-1.5, -12.5), and its y-intercept is (0, -8). The x-intercepts of the parabola are

OpenStudy (anonymous):

we can use : y=m (x-h)+ k Vertex= (h,k) or m= y2-y1/ x2x1 y-y1= m (x-x1)

OpenStudy (anonymous):

im lost can you put the numbers in for me